# GlobalTomo: A global dataset for physics-ML seismic wavefield modeling and FWI

### v3

@@ 4 Conclusion
we introduce GlobalTomo, a 3D global full waveform dataset for physics-ML research in forward modeling and FWI. Our analysis showcases the strengths and limitations of modern ML methods’ capabilities in seismic wavefield modeling, providing potential tools and methods for forward modeling and inversion. Through this cross-disciplinary endeavor, we anticipate that the convergence of advanced machine learning techniques and traditional geophysical methods will accelerate discoveries in earth science, promoting a deeper understanding of the earth’s interior dynamics.
@@ C.2 Three Dataset Tiers
This set of models is designed for the ultimate real-earth applications. We utilize the isotropic PREM [12] background model and a global mesh with a period accurate to 30 seconds that encompasses various earth’s layers including the crust, mantle, outer core, and inner core, each characterized by their unique seismic attributes. The Real Earth tier has a stress-free boundary condition. The 3D velocity perturbations are imposed on the earth’s internal 23 discontinuities, resulting in a total of 5427 free parameters up to degree 8 across the whole earth.
@@ Appendix E Comparison with Other Related Datasets
We contrasts the GlobalTomo dataset with prior works across several key aspects in Table A1. The comparison highlights the differences in scale, data utilization, parameterization, and simulation methods. GlobalTomo is first designed for global-scale inversion, incorporating the entire wavefield and receiver data, utilizing spherical harmonics for structural parameterization, and employing spectral element methods (SEM) for efficient simulation. In contrast, prior works (e.g., OpenFWI [10], E-FWI [14]) typically focus on local-scale applications, use only receiver data, rely on mesh representation, and often employ finite difference methods (FDM) for simulation. DiTing [76] and SeisBench [72] are particularly designed for phase picking, denoising, and earthquake detection instead of earth tomography.
@@ F.1 Evaluation Metrics
Relative $\text{Relative }L_{2}\text{ Loss}=\sqrt{\frac{\sum_{i=1}^{n}(y_{i}-\hat{y}_{i})^{2}}{\sum_{i=1}^{n}(y_{i}-\overline{y})^{2}}}$ (A22)
@@ F.2 Efficient DeepONet Design
However, this is memory-consuming since the velocity structures are duplicated for each point. Indeed, one velocity structure can be paired with a group of spatial and temporal points $P,T$ simultaneously. Inspired by [40], we employ a more efficient DeepONet that can predict the whole 3D outputs just using one forward of branch net. The calculation of the output function now is given by:
@@ G.1 Comparison of Modern ML-based Inversion
In comparing the inversion strategies of PINN and operator learning, we observe distinct paths of optimization. The PINN-based approach seeks to optimize the model parameters $c$ in conjunction with the forward modeling function $\Phi$ [51]. Thus, the $c$ and $\Phi$ are closely coupled and the inversion of different structures requires training from scratch. This method limits the model’s ability to generalize across different velocity structures.
In contrast, the operator learning approach addresses this limitation by initially learning a universal operator to model forward problems across all $c$ [74]. Subsequently, it fine-tunes the model parameters $c$ for each specific case using the pre-trained $\Phi$. Once trained, the neural operator can achieve inversion on any structures in inference time.
@@ G.3 Connection with Other Fields
Both seismic inversion and intuitive physics deal with inverse problems—inferring causes from effects. Seismic inversion back-calculates to infer the earth’s interior from surface measurements [65]. Similarly, intuitive physics tries to deduce the cognitive rules or mental models that people use to predict physical outcomes from their observed behavior in experimental settings [28].
Regarding methodologies, both fields heavily utilize computational models. In seismic studies, numerical methods and statistical models predict how geological features influence seismic waveforms [63, 66]. Similarly, in intuitive physics, physics engines simulate how humans might predict physical interactions in their surroundings [2, 34]. Machine learning, particularly neural networks, plays a significant role in both areas. In seismic inversion, neural networks can be trained to recognize patterns in seismic data that correlate with specific subsurface structures [74]. In intuitive physics, neural networks model human prediction and reasoning processes about physical laws, learning to mimic human judgment [33].
Uncertainty management is crucial in both domains. In seismic inversion, uncertainties arise from limited sensor coverage and noise in data. In intuitive physics, uncertainties stem from the variability in human perception and cognitive biases. Both fields develop methodologies to handle these uncertainties, often through probabilistic models [65, 2].
@@ I.3 A Checkboard Test
To validate the robustness of our inversion methods, we conducted a checkerboard test as a standard diagnostic tool in geophysical tomography [52, 65]. The test involves simulating a known subsurface model with alternating high and low-velocity regions arranged in a checkerboard pattern, which allows us to evaluate the algorithm’s ability to recover fine-scale variations in the subsurface. We evaluate the checkboard inversion using our pre-trained InversionMLP. The inversion results demonstrated a high correlation of 0.89 with the ground truth structures, with the algorithm successfully resolving the checkerboard structure and accurately recovering the modeled velocity contrasts; see Figure A21 for visualization. This robust performance indicates that our approach is capable of handling complex geophysical inversion tasks and offers significant potential for real-world applications in seismic tomography and other geophysical imaging techniques.

### v4

@@ 4 Conclusion
In this work, we introduce GlobalTomo, a comprehensive 3D global full waveform dataset for physics-ML research in forward modeling and FWI. By synthesizing high-fidelity seismic data across a range of scales and complexities with efficient spherical parameterization, GlobalTomo facilitates the advanced training and evaluation of ML models in the realm of seismic tomography. Our extensive evaluations confirm that ML-based approaches, supported by GlobalTomo, substantially enhance the efficiency and scalability of seismic wavefield modeling and FWI processes. Moreover, this dataset offers new possibilities for integrating machine learning with physics-based modeling. Through this cross-disciplinary endeavor, we anticipate that the convergence of advanced machine learning techniques and traditional geophysical methods will accelerate discoveries in earth science, promoting a deeper understanding of the earth’s interior dynamics and supporting critical applications such as resource discovery and hazard assessment.
Impact statements
A deeper understanding of the earth’s internal structure brings numerous societal benefits, including improved disaster preparedness and mitigation, safer urban planning and infrastructure development, more effective and sustainable management of natural resources, and environmental preservation. This research also fosters global scientific and technological collaboration, further amplifying its positive societal impacts.
Acknowledgments and Disclosure of Funding
The authors would like to thank Miss Chen Zhen (BIGAI) for making the nice figures, Prof. Yizhou Wang (Peking University) for supporting the project, and NVIDIA for their valuable technical support for accelerated computing and artificial intelligence. S. Li, Z. Mu, S. Xin, and Y. Zhu are supported in part by the National Science and Technology Major Project (2022ZD0114900), the National Natural Science Foundation of China (62376009), the Beijing Nova Program, the State Key Lab of General AI at Peking University, the PKU-BingJi Joint Laboratory for Artificial Intelligence, and the National Comprehensive Experimental Base for Governance of Intelligent Society, Wuhan East Lake High-Tech Development Zone. Z. Li and X. Song are supported in part by the National Key R&D Program of China (2022YFF0800601), China Postdoctoral Science Foundation (2023M740043), and the Emerging Engineering Interdisciplinary–Young Scholars Project, Fundamental Research Funds for the Central Universities. K. Leng is supported in part by the EPSRC grant, Blueprinting for AI for Science at Exascale (BASE-II, EP/X019918/1).
@@ C.2 Three Dataset Tiers
This set of models is designed for the ultimate real-earth applications. We utilize the isotropic PREM (Dziewonski and Anderson, , 1981) background model and a global mesh with a period accurate to 30 seconds that encompasses various earth’s layers including the crust, mantle, outer core, and inner core, each characterized by their unique seismic attributes. The Real Earth tier has a stress-free boundary condition. The 3D velocity perturbations are imposed on the earth’s internal 23 discontinuities, resulting in a total of 5427 free parameters up to degree 8 across the whole earth.
@@ Appendix E Comparison with Other Related Datasets
We contrasts the GlobalTomo dataset with prior works across several key aspects in Table A1. The comparison highlights the differences in scale, data utilization, parameterization, and simulation methods. GlobalTomo is first designed for global-scale inversion, incorporating the entire wavefield and receiver data, utilizing spherical harmonics for structural parameterization, and employing spectral element methods (SEM) for efficient simulation. In contrast, prior works (e.g., OpenFWI (Deng et al., , 2022), E-FWI (Feng et al., , 2024)) typically focus on local-scale applications, use only receiver data, rely on mesh representation, and often employ finite difference methods (FDM) for simulation. DiTing (Zhao et al., , 2023) and SeisBench (Woollam et al., , 2022) are particularly designed for phase picking, denoising, and earthquake detection instead of earth tomography.
@@ F.1 Evaluation Metrics
$\text{Relative }L_{2}\text{ Loss}=\sqrt{\frac{\sum_{i=1}^{n}(y_{i}-\hat{y}_{i})^{2}}{\sum_{i=1}^{n}(y_{i}-\overline{y})^{2}}}$ (A22)
@@ F.2 Efficient DeepONet Design
However, this is memory-consuming since the velocity structures are duplicated for each point. Indeed, one velocity structure can be paired with a group of spatial and temporal points $P,T$ simultaneously. Inspired by (Lu et al., , 2022), we employ a more efficient DeepONet that can predict the whole 3D outputs just using one forward of branch net. The calculation of the output function now is given by:
@@ G.1 Comparison of Modern ML-based Inversion
In comparing the inversion strategies of PINN and operator learning, we observe distinct paths of optimization. The PINN-based approach seeks to optimize the model parameters $c$ in conjunction with the forward modeling function $\Phi$ (Rasht-Behesht et al., , 2022). Thus, the $c$ and $\Phi$ are closely coupled and the inversion of different structures requires training from scratch. This method limits the model’s ability to generalize across different velocity structures.
In contrast, the operator learning approach addresses this limitation by initially learning a universal operator to model forward problems across all $c$ (Yang et al., , 2023). Subsequently, it fine-tunes the model parameters $c$ for each specific case using the pre-trained $\Phi$. Once trained, the neural operator can achieve inversion on any structures in inference time.
@@ G.3 Connection with Other Fields
Both seismic inversion and intuitive physics deal with inverse problems—inferring causes from effects. Seismic inversion back-calculates to infer the earth’s interior from surface measurements (Tromp, , 2020). Similarly, intuitive physics tries to deduce the cognitive rules or mental models that people use to predict physical outcomes from their observed behavior in experimental settings (Kubricht et al., , 2017).
Regarding methodologies, both fields heavily utilize computational models. In seismic studies, numerical methods and statistical models predict how geological features influence seismic waveforms (Tarantola, , 1984; Tromp et al., , 2005). Similarly, in intuitive physics, physics engines simulate how humans might predict physical interactions in their surroundings (Battaglia et al., , 2013; Li et al., , 2023). Machine learning, particularly neural networks, plays a significant role in both areas. In seismic inversion, neural networks can be trained to recognize patterns in seismic data that correlate with specific subsurface structures (Yang et al., , 2023). In intuitive physics, neural networks model human prediction and reasoning processes about physical laws, learning to mimic human judgment (Li et al., 2022a, ).
Uncertainty management is crucial in both domains. In seismic inversion, uncertainties arise from limited sensor coverage and noise in data. In intuitive physics, uncertainties stem from the variability in human perception and cognitive biases. Both fields develop methodologies to handle these uncertainties, often through probabilistic models (Tromp, , 2020; Battaglia et al., , 2013).
@@ I.3 A Checkboard Test
To validate the robustness of our inversion methods, we conducted a checkerboard test as a standard diagnostic tool in geophysical tomography (Rawlinson et al., , 2014; Tromp, , 2020). The test involves simulating a known subsurface model with alternating high and low-velocity regions arranged in a checkerboard pattern, which allows us to evaluate the algorithm’s ability to recover fine-scale variations in the subsurface. We evaluate the checkboard inversion using our pre-trained InversionMLP. The inversion results demonstrated a high correlation of 0.89 with the ground truth structures, with the algorithm successfully resolving the checkerboard structure and accurately recovering the modeled velocity contrasts; see Figure A21 for visualization. This robust performance indicates that our approach is capable of handling complex geophysical inversion tasks and offers significant potential for real-world applications in seismic tomography and other geophysical imaging techniques.